OpenStudy (anonymous):

Use the indicated substitution to evaluate the integral : The Integral of (1-cos(t/2))^2 x (sin(t/2)) dt u= 1-cos(t/2)

OpenStudy (slaaibak):

du = sin(t/2)/2 dt so integral becomes (u^2) 2du = 2u^3 /3 + C just sub u= 1-cos(t/2) back

OpenStudy (anonymous):

Where does the division ( the sin(t/2) / 2 come from.

OpenStudy (slaaibak):

differentiating u= 1-cos(t/2)

OpenStudy (anonymous):

I think you are thinking of something else. You would then have to solve for dt first.... i.e. dt= 2du/sin(t/2) and then I would plug it back in. But thanks.

OpenStudy (slaaibak):

I'm correct. I've done this plenty of times.

OpenStudy (slaaibak):

I'll fully type it out with latex if you don't know what I'm saying.

OpenStudy (anonymous):

Hahaha you got the wrong answer though. Its chill. BC Calc sucks .

OpenStudy (slaaibak):

\[\int\limits (1 - \cos ({t \over 2}))^2 \times \sin({t \over 2}) dt\] now, firstly: \[u = 1 - \cos({t \over 2}) \rightarrow du = {1 \over 2}\sin({t \over 2}) dt \rightarrow \sin({t \over 2}) dt = 2du\] now, replace sin(t/2)dt with 2 du and 1- cos(t/2) with u the integral becomes: \[\int\limits 2u^2 du = {2 \over 3} u^3 + C = {2 \over 3}(1-\cos({t \over 2}))^3 + C\]

OpenStudy (slaaibak):

This answer is correct.

OpenStudy (anonymous):

k that made a lot more sense.

OpenStudy (anonymous):

Actually you need to pull it to .....^3/2 i.e. anti derivative formula of: \[(X^(n+1))/ (n+1) \]

OpenStudy (slaaibak):

I did that.

OpenStudy (anonymous):

You left it as cubed rather than 3/2

OpenStudy (slaaibak):

why on earth would it be 3/2?

OpenStudy (slaaibak):

\[\int\limits u^2 = {u^3 \over 3} +C\]

OpenStudy (anonymous):

Mentor teaching the mentee huh. When you take the Anti-Derivative of \[\sqrt{\tan x}\] The formula that I just showed you can be applied. So \[\sqrt{\tan x}\] is technically \[\tan^.5\] Correct? Putting it into the equation you would get that result.

OpenStudy (anonymous):

Of 2/3 tan ^ 3/2

OpenStudy (slaaibak):

You're teaching me how to laugh at you, you're not teaching me any math. THERES NO SQUARE ROOT IN YOUR QUESTION.

OpenStudy (anonymous):

whoopsy looking at the wrong equations. Don't get butt hurt guy.

OpenStudy (slaaibak):

don't try and act smart.

OpenStudy (anonymous):

Soooooooo Butt hurt. My GPA laughs at you.

OpenStudy (slaaibak):

Just try and get through high school.

OpenStudy (slaaibak):

oh, and by the way, if you can't do an integral as simple as this, I highly doubt your GPA is even able to laugh.

OpenStudy (anonymous):

4.0 Unweighted. And its sad to see you just learning this in College but I don't judge.

OpenStudy (slaaibak):

I didn't learn this in college. I learned this at school. The fact that I can do this while I'm in university says nothing about when I learnt this. I'm not a math major either.